OpenStudy (anonymous):

Find two consecutive integers whose sum is 67. Which of the following equations could be used to solve the problem? 2x = 67 2x + 1 = 67 2x + 2 = 67 x2 + 1= 67

OpenStudy (anonymous):

2x + 1 = 67 because we can say that X is the first integer, than the second integer would be X+1 since they are two consecutive integers. so X + (X+1) = 67 --> 2X + 1 = 67 And in case you need it, the two integers would be 33 and 34 :)